(Chapter 7) You analyze the results of a study using a t-test for dependent means. There were 45 participants, each tested before and after. The t-score for your result was 2.89 which was significant at the .01 level.   How would you report this result in a journal article (just give symbols--similar to depiction below)? Answer in terms of t(x)=2.89 p<0.1
Answer #1

P(t > 2.89 ) at df = n-1 is 45-1 = 44

= 1- t.dist(2.89 , 44, true)

= .00298 or .003

So, this is less than p = .010

Ans hence we conclude that the result is statistically significant., we reject Ho.